This is one question with different parts, please help. Thank you
Find the vertex, line of symmetry, the maximum or minimum value of the quadratic function
f(x)=-2x^2+2x+3
What is the x-coordinate of the vertex?
What is the y-coordinate of the vertex?
What is the equation of the line of symmetry?
The max/min of f(x) is?
The value, f(1/2)=7/2 is a maximum or a minimum?

Answer by stanbon(75887)

Find the vertex, line of symmetry, the maximum or minimum value of the quadratic function
f(x)=-2x^2+2x+3
Complete the square to see the vertex form of the quadratic:
-2x^2 + 2x + ? = y - 3 + ?
-2(x^2 -x + (1/4)) = y - 3 - 2(1/4)
-2(x-(1/2))^2 = y - 3.5
-------------------------
What is the x-coordinate of the vertex?
x = 1/2
-----------------
What is the y-coordinate of the vertex?
y = 3.5
---------------------------
What is the equation of the line of symmetry?
x = 1/2
---------------------------------
f(x)=-2x^2+2x+3
---------------------
The max/min of f(x) is?
Min or max occurs when x = -b/2a = -2/(2*-2) = 1/2
----------------------------------------
The value, f(1/2)=7/2 is a maximum or a minimum?
It is a max because the coefficient of x^2 is negative.
===========================================================
Cheers,
Stan H.